Grade 12th passAlgebra

How many ways are there to arrange the letter in the word Sachin with the vowels in alphabetical order

Atul Tiwari
9 Years agoGrade 12th pass

1 Answer

Pratyush Ranjan Roul
9 Years ago
As “a , i ” are the two vowels
So, there are six conditions :-
1)  a i _ _ _ _
​2) a _ i _ _ _
​3) a _ _ i _ _
​4) a _ _ _ i _
​5) a _ _ _ _ i
In the dashes the letters “ S, c, h, n” can be filled 
As there are no repetitions 
So,
the no. of ways in all cases are 4!  i.e. = 24
So, as there are 6 separate ways of doing so
Therefore we add it up =6×24 = 144
But, the i can be replaced by a
So,
Total no. of ways = 2×144=288